OpenStudy (anonymous):

tracy needs to find the length and width of the equipment storage room where she works . the room is 3 meters longer than it is wide its perimeter is 18 meters. find the width in meters

OpenStudy (nurali):

Let the width of the room = x m then the length = (x+3) m Perimeter of the room = 2(Length+Width) 2(x+3+x) = 18 4x+6 = 18 4x = 18-6 = 12 x = 12/4 = 3 Therefore the width of the room = 3 meters